OpenStudy (anonymous):

Find the circumference of the circle. Use p » 3.14. A. 3 m B. 18.84 m C. 6 m

OpenStudy (anonymous):



OpenStudy (anonymous):

6m ????

OpenStudy (anonymous):

you can use the formula circumference = 2 pi r

OpenStudy (anonymous):

\[2\pi r\]

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

substitute the value of r in the formula

OpenStudy (anonymous):

3*3.14=942

OpenStudy (anonymous):

i dont know what to do

OpenStudy (anonymous):

2*3*3.14=B

OpenStudy (anonymous):

yaaa..